Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $3,000. It seems
Answer #1

1a) Predetermine overhead rate = 911400/651000 = 140% of labor cost

1b) Overhead applied

Fabricating Machining Assembly Total
Overhead applied 6200*1.4 = 8680 600*1.4 = 840 7900*1.4 = 11060 20580

2a) Departmental overhead rate :

Fabricating = 379750/217000 = 175% of labor cost

Machining = 434000/108500 = 400% of labor cost

Assembly = 97650/325500 = 30% of labor cost

2b) Overhead applied

Fabricating Machining Assembly Total
Overhead applied 6200*1.75 = 10850 600*4 = 2400 7900*30% = 2370 15620

3a) Bid price = (8100+14700+20580)*1.5 = 65070

3b) Bid price = (8100+14700+15620)*1.5 = $57630